ABC Limited is looking at expanding its business and wants to invest in a new plant to boost its production capacity. The plant has a life of three years. The details are as follows.

 The plant would depreciate in three years from the acquiring cost of Rs. 4,20,000 to zero in three years. There would be no salvage value at the end of three years. The depreciation would be on a straight line basis.

 The additional revenue from the plant would be ~ Rs. 6,00,000 in year 1, Rs. 7,00,000 in year 2 and 3.

 The input cost (raw material) is expected to be ~ Rs. 3,00,000 for year 1 and 2 and Rs. 4,00,000 in year 3.

Assuming a tax rate of 30% and a discount rate of 20%, you are required to

1. Arrive at the expected annual cash flows (after-tax)


2. Compute the net present value of the investment and determine if the investment is financially viable?


1
Expert's answer
2021-09-21T05:03:17-0400



Initial investment=420 000

Income receipts=600 000, 700 000, 700 000.

"Depreciation=\frac{1}{n}\times 420 000=\frac{1}{3}\times 420 000=140 000"

n=3 years

Expenses=300 000, 400 000, 400 000.

Cash flow is equal to depreciation and minus expenses:

600 000-140 000-300 000=160 000 for the first year and so on.

Net cash flow is a shareable flow minus tax:

"160 000-0.3\times160 000=112 000" for the first year and so on.

Discounted cash flow is net cash flow divided by the discount rate:

"\times{112 000}{(1+0.2)^n}=\frac{112 000}{1.2}=93 333.33" for the first year and so on.

n=1,2.3 by year, respectively

Net discounted income is the sum of the initial investment with a minus sign adding the discounted cash flows:

NPV=-420 000+93 333.33+126 388.89+64 814.81=-135 463

If the NPV is greater than zero, then the project can be accepted, if the NPV is less than zero, then the project should be rejected. A negative NPV value indicates that the project is not attractive.